OpenStudy (anonymous):

Suppose y varies jointly as x and z. Find y when x=2and z =27, if y = 192 when x=8 and z= 6.

OpenStudy (anonymous):

@moha_10 do u kno how to do this?

OpenStudy (anonymous):

It is a direct variation: \[\large y = k \times (xz)\] where k is any constant.. Now you are given with y = 192 x = 8 and z = 6 can you find k from here..

OpenStudy (anonymous):

@janessia atleast you should reply in case you are not getting...

OpenStudy (anonymous):

im sorry i was helping my mom gosh!

OpenStudy (anonymous):

hold on i maybe can find k let me try

OpenStudy (anonymous):

does k=4? @waterineyes

OpenStudy (anonymous):

Yes you are right.. Again use the same formula: Now you have k=4 x = 2 an d z = 27 find y..

OpenStudy (anonymous):

216?

OpenStudy (anonymous):

Yes this is the value of y.. Well done...

OpenStudy (anonymous):

yayy! i have another question like this If y varies inversely as x and y=2 when x =25 find x when y=40. @waterineyes

OpenStudy (anonymous):

In case of inverse proportion: k remains in the numerator but x or whatever is given goes to denominator like this: \[\large \color{green}{y = \frac{k}{x}}\] First of all find k by putting x = 25 and y = 2.. After that when you got k using the same formula put y = 40 and find x from there..

OpenStudy (anonymous):

ohk hold on

OpenStudy (anonymous):

ohk i got a big number i got 2000

OpenStudy (anonymous):

what is k??

OpenStudy (anonymous):

i got 50 for k

OpenStudy (anonymous):

Yes now you have y = 40 and k = 50 solve for x..

OpenStudy (anonymous):

why are you multiplying them:???

OpenStudy (anonymous):

yes when u look for x u multiply right? cuz its being divided?

OpenStudy (anonymous):

\[y = \frac{k}{x}\] \[x = \frac{k}{y}\]

OpenStudy (anonymous):

ohhhhh

OpenStudy (anonymous):

i got 1.25

OpenStudy (anonymous):

Hurry up tell me It is 3:21 in the morning here.. I have to shut my pc down.. Yes it is 5/4 = 1.25..

OpenStudy (anonymous):

Good Night janessia.. Well done....

OpenStudy (anonymous):

thnk u

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

It is okay no need to be sorry.. Welcome dear..

OpenStudy (anonymous):

:)